Probability

Get insights from 304 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability

  • 304

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 4

    Followers

Have a question about Probability? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

39. Let, E1: event person has disease

E2: event person has no disease

A: be event that blood test is positive

As E1 and E2 are events which are complementary to each other,

Then, P(E1)+P(E2)=1

P(E2)=1−P(E1)

P(E1)=0.1%=0.1100=0.001

∴ P(E2)=1−0.001

=0.999

P(A/E1)=P (result is positive given that person has disease) =99%=0.99

P(A/E2)=P (result is positiv

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

38. Let, E1 :event that knows the answer

E2 :event that he guesses the answer

P(E1)=34 and P(E2)=14

Let A be the event that the answer is correct

Also, P(A/E1)=P (correct answer given that he knows) =1

P(A/E2)=P (correct answer given that he guess) =14

Now, probability that he knows the an

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

37. Let, E1 :event that the student in hostel

E2 : event that the student is day scholar

A : be the event of the chosen student get grade A

P(E1)=60%=60100=0.6

P(E2)=40%=40100=0.4

P(A/E1)=P (student getting an A grade is hostler) =30% =30100=0.3

P(A/E2)=P ( student getting an A grade is a day scholar) =20%=20100=0.2

Then, by Ba

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

36. Let E1 be the event of selecting first bag

E2 be the event of selecting second bag

∴P(E1)=P(E2)=12

Let A be the event of getting a red ball

P(A/E1)=P (drawing a red ball from first bag)

=48=12

P(A/E2)=P (drawing a red ball from second bag)

=28=14

Then, probability of drawing a ball from the first bag which is red, is

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

35. Number of balls contain in win =5 red

=5 black

Let the red ball be drawn in the first attempt

∴ P (drawing a red ball) =510=12

By question, if added two red balls to the win, then 7 red balls and 5 black balls contain.

P (drawing a red ball) =712

Let a black ball be

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

34. P (A′B′)= [1−P (A)] [1−P (B)]

⇒P (A′∩B′)=1−P (A)−P (B)+P (A)P (B)

⇒− [P (A)+P (B)−P (A∩B)]=−P (A)−P (B)+P (A) (B)

⇒−P (A)−P (B)+P (A∩B)=−P (A)−P (B)+P (A)P (B)

⇒P (A∩B)=P (A).P (B)

Therefore, it shows A and B are independant

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

33. The outcome of rolling 2 dice is 36 . The only prime number is 2 .

Let,  E= event of getting an even prime number on each die

E= { (2, 2)} therefore

P (E)=136

Therefore, the correct answer is  (D) 136

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

32. Let,

P(H)= denote the student who read Hindi newspaper

P(E)= denote the student who read English newspaper

Then, P(H)=60%=610=35

P(E)=40%=410=25

i. P(H∩E)=20%=210=15

⇒P(H∪E)′=1−[P(H)+P(E)−P(H∩E)]

=1−[35+25−15]

=1−3+2−15

=1−45=15

ii. Probability of randomly chosen student that reads English newspaper, if she reads Hindi newspaper, is given by

P(E/H)=P(E∩H)P(H)=1535=13

iii. Probability that randomly chosen student reads Hindi newspaper

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

31. i. In a deck of 52 cards,

13 are spades and 4 are aces.

∴P(E)=P (the card drawn is spade) =1352=14

∴P(F)=P (the card drawn is an ace) =452=113

∴P(E∩F)=P (the card drawn is spade and an ace) =152

P(E).P(F)=14×113=152=P(E∩F)

⇒P(E).P(F)=P(E∩F)


E and F are independent.

ii. In a deck of 52 cards,

26 cards are black and 4 a

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

30. Let, P(A)= probability of solving problem by A=12

P(B)= probability of solving problem by B=13

Since, the problem is solved independently by A and B , then,

⇒P(A∩B)=P(A).P(B)

⇒P(A∩B)=12×13

=16

P(A′)=1−P(A)=1−12=12

P(B′)=1−P(B)=1−13=23

i. Probability of the problem is solved

⇒P(A∪B)=P(A)+P(B)−P(A∩B)

⇒P(A∪B)=12+13−16

=3+2−16

=46=23

ii. Probability that exactly one of the solved problem is given by P(A)P(B′)+P(B)P(A′)=12×23+12×13

=26+16=36=12

Thumbs Up IconUpvote Thumbs Down Icon